# Mean-Field Sampling for Cooperative Multi-Agent Reinforcement Learning

### v1

@@ Appendix F Using the Performance Difference Lemma
$\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i,\Delta^{i}},\pi^{*}(s)_{g},\pi^{*}(s)_{i,\Delta^{i}})\right|$
$\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}\left(1-\mathbb{E}_{s^{\prime}\sim\mathcal{D}}\mathcal{I}(s^{\prime},k,\delta)\right))$
$\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
$\quad=\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
For the first term in the first inequality, we use $\mathbb{E}[X\mathbbm{1}\{X\leq c\}]\leq c$. For the second term, we trivially bound $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\widehat{Q}_{k}^{*}(s_{g},s_{i,\Delta^{i}},\pi^{*}(s)_{g},\pi^{*}(s)_{i,\Delta^{i}})$ by the maximum value $Q^{*}$ can take, which is $\frac{\widetilde{r}}{1-\gamma}$ by Lemma B.5. In the second inequality, we use the fact that the expectation of an indicator function is the conditional probability of the underlying event. The second inequality follows from Lemma F.5.
@@
$\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s,\pi^{*}(s))-\widehat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\right|$
$\quad\quad\quad\quad\quad\quad\quad\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
@@
Corollary F.6.1.
Changing the notation of the policy to ${\pi}_{k,m}^{\mathrm{est}}$ in the proofs of Lemmas F.6 and F.5,such that $a\sim\widetilde{\pi}_{k,m}^{\mathrm{est}}$ then yields that:
$\mathbb{E}_{s\sim\mathcal{D}}\left[\sum_{\Delta\in{[n]\choose k}}\sum_{\Delta^{1}\in{[n]\setminus 1\choose k-1}}\sum_{\Delta^{2}\in{[n]\setminus 2\choose k-1}}\dots\sum_{\Delta^{n}\in{[n]\setminus n\choose k-1}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\left|Q^{*}(s,a)-\widehat{Q}^{*}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},a_{g},a_{i,\Delta^{i}\}})\right|\right]$
$\quad\quad\quad\quad\leq\frac{\widetilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\widetilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
Lemma F.7 (Optimizing Parameters).
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right)$
@@
Setting $\delta=\frac{(1-\gamma)^{2}}{2\widetilde{r}\varepsilon|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}}$ in the bound for the optimality gap in Theorem F.3 gives:
$V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}(s_{0})$ $\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+2\epsilon_{k,m}+\frac{\widetilde{2r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
$=\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{4\widetilde{r}\varepsilon|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}}{(1-\gamma)^{2}}}+\frac{1}{\varepsilon}+2\epsilon_{k,m}$
@@
$V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}^{\mathrm{est}}}(s_{0})\leq\frac{\widetilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{40\widetilde{r}|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|+\frac{1}{2}}}{(1-\gamma)^{2}}}+\frac{1}{10\sqrt{k}}+2\epsilon_{k,m}$
Finally, setting $\epsilon_{k,m}\leq\widetilde{O}(\frac{1}{\sqrt{k}})$ from Lemma F.9 yields that
$V^{\pi^{*}}(s_{0})-V^{\widetilde{\pi}_{k,m}}(s_{0})\leq\widetilde{O}\left(\frac{1}{\sqrt{k}}\right),$
@@
F.1 Proof of Lemma 3.2: Bounding the Bellman Error
Theorem F.8 (Theorem 2 of Li et al. (2022)).
@@
$\|\widehat{\mathcal{T}}_{k,m}\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{\mathcal{T}}_{k}\widehat{Q}_{k}^{*}\|_{\infty}=\|\widehat{Q}_{k,m}^{\mathrm{est}}-\widehat{Q}_{k}^{*}\|_{\infty}\leq\epsilon_{k,m},$
@@
$m=\frac{c_{0}\cdot\widetilde{r}\cdot t_{\text{cover}}}{(1-\gamma)^{5}\epsilon_{k,m}^{2}}\log^{2}(|\mathcal{S}||\mathcal{A}|)\log\left(\frac{1}{(1-\gamma)^{2}}\right),$ (33)
where $t_{\text{cover}}$ stands for the cover time, and is the time taken for the trajectory to visit all state-action pairs at least once. Formally,
@@
Lemma F.9.
The SUB-SAMPLE-MFQ: Learning algorithm runs in time $\widetilde{O}(|\mathcal{S}_{g}|^{2}|\mathcal{A}_{g}|^{2}|\mathcal{A}_{l}|^{2}|\mathcal{S}_{l}|^{2}k^{2.5+2|\mathcal{S}_{l}||\mathcal{A}_{l}|}\widetilde{r})$, while attaining a Bellman noise $\epsilon_{k,m}$ of $\widetilde{O}(1/\sqrt{k})$.
@@
Rearranging Eq. 33, one has
$\epsilon_{k,m}\leq\frac{k\sqrt{2\widetilde{r}t_{\text{cover}}}}{(1-\gamma)^{2.5}\sqrt{m}}\log\left(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|\right)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$ (35)
@@
Substituting this in Eq. 35 gives
$\epsilon_{k,m}\leq\frac{k\sqrt{2\widetilde{r}|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|}}}{(1-\gamma)^{2.5}\sqrt{m}}\log\left(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|\right)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$ (36)
Therefore, setting
$m=\frac{2\widetilde{r}|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{2.5+|\mathcal{S}_{l}||\mathcal{A}_{l}|}}{(1-\gamma)^{5}}\log(|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$ (37)
attains a Bellman error of $\epsilon_{k,m}\leq\widetilde{O}(1/\sqrt{k})$. Finally, the runtime of our learning algorithm is
$O(m|\mathcal{S}_{g}||\mathcal{S}_{l}||\mathcal{A}_{g}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|})=\widetilde{O}(|\mathcal{S}_{g}|^{2}|\mathcal{A}_{g}|^{2}|\mathcal{A}_{l}|^{2}|\mathcal{S}_{l}|^{2}k^{2.5+2|\mathcal{S}_{l}||\mathcal{A}_{l}|}\widetilde{r}),$
@@
Appendix G Generalize to Stochastic Rewards

### v3

@@ Appendix G Using the Performance Difference Lemma
$\mathbb{E}_{s\sim\mathcal{D}}\left|Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i,\Delta^{i}},\pi^{*}(s)_{g},\pi^{*}(s)_{i,\Delta^{i}})\right|$
$\quad\leq\frac{\tilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\tilde{r}}{1-\gamma}\left(1-\mathbb{E}_{s^{\prime}\sim\mathcal{D}}\mathcal{I}(s^{\prime},k,\delta)\right))$
$\quad\leq\frac{\tilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\tilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
$\quad=\frac{\tilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\tilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
For the first term in the first inequality, we use $\mathbb{E}[X\mathbbm{1}\{X\leq c\}]\leq c$. For the second term, we trivially bound $Q^{*}(s^{\prime},\pi^{*}(s^{\prime}))-\hat{Q}_{k}^{*}(s_{g},s_{i,\Delta^{i}},\pi^{*}(s)_{g},\pi^{*}(s)_{i,\Delta^{i}})$ by the maximum value $Q^{*}$ can take, which is $\frac{\tilde{r}}{1-\gamma}$ by Lemma C.4. In the second inequality, we use the fact that the expectation of an indicator function is the conditional probability of the underlying event. The second inequality follows from Lemma G.5.
@@
$\!\!\sum_{\Delta\in{[n]\choose k}}\!\sum_{\begin{subarray}{c}\Delta^{1},\dots,\Delta^{n}:\\ \Delta^{i}\in{[n]\setminus i\choose k-1}\end{subarray}}\!\frac{1/n}{{n\choose k}{n-1\choose k-1}^{n}}\!\sum_{i\in[n]}\mathbb{E}_{s\sim\mathcal{D}}\bigg{|}Q^{*}(s,\pi^{*}(s))\!-\!\hat{Q}_{k,m}^{\mathrm{est}}(s_{g},s_{i},F_{s_{\Delta^{i}}},\pi^{*}(s)_{g},\{\pi^{*}(s)_{j}\}_{j\in\{i,\Delta^{i}\}})\bigg{|}$
$\quad\quad\quad\quad\quad\quad\quad\quad\leq\frac{\tilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\tilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
@@
Corollary G.7.
Changing the notation of the policy to ${\pi}_{k,m}^{\mathrm{est}}$ in the proofs of Lemmas G.6 and G.5, such that $a\sim\tilde{\pi}_{k,m}^{\mathrm{est}}$ then yields that:
$\mathbb{E}_{s\sim\mathcal{D}}\left[\sum_{\Delta\in{[n]\choose k}}\sum_{\begin{subarray}{c}\Delta^{1},\dots,\Delta^{n}:\Delta^{i}\in\in{[n]\setminus i\choose k-1}\end{subarray}}\frac{1}{{n\choose k}}\frac{1}{{n-1\choose k-1}^{n}}\sum_{i\in[n]}\frac{1}{n}\left|Q^{*}(s,a)-\hat{Q}^{*}_{k,m}(s_{g},s_{i},F_{s_{\Delta^{i}}},a_{g},a_{i,\Delta^{i}\}})\right|\right]$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\leq\frac{\tilde{r}}{1-\gamma}\sqrt{\frac{n-k+1}{8nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+\epsilon_{k,m}+\frac{\tilde{r}}{1-\gamma}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
Lemma G.8 (Optimizing Parameters).
With probability at least $1-\frac{1}{100e^{k}}$,
$V^{\pi^{*}}(s_{0})-V^{{\pi}_{k,m}}(s_{0})\leq\tilde{O}\left(\frac{1}{\sqrt{k}}\right)$
@@
Setting $\delta=\frac{(1-\gamma)^{2}}{2\tilde{r}\varepsilon|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}}$ in the bound for the optimality gap in Theorem G.3 gives:
$V^{\pi^{*}}(s_{0})-V^{\tilde{\pi}_{k,m}^{\mathrm{est}}}(s_{0})$ $\leq\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{2|\mathcal{S}_{l}||\mathcal{A}_{l}|}{\delta}}+2\epsilon_{k,m}+\frac{\tilde{2r}}{(1-\gamma)^{2}}|\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}\delta$
$=\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{4\tilde{r}\varepsilon|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|}}{(1-\gamma)^{2}}}+\frac{1}{\varepsilon}+2\epsilon_{k,m}$
@@
$V^{\pi^{*}}(s_{0})-V^{\tilde{\pi}_{k,m}^{\mathrm{est}}}(s_{0})\leq\frac{\tilde{r}}{(1-\gamma)^{2}}\sqrt{\frac{n-k+1}{2nk}}\sqrt{\ln\frac{40\tilde{r}|\mathcal{S}_{l}||\mathcal{A}_{l}||\mathcal{A}_{g}|k^{|\mathcal{A}_{l}|+\frac{1}{2}}}{(1-\gamma)^{2}}}+\frac{1}{10\sqrt{k}}+2\epsilon_{k,m}$
Finally, using the probabilistic bound of $\epsilon_{k,m}\leq\tilde{O}(\frac{1}{\sqrt{k}})$ from Lemma G.10 yields that with probability at least $1-\frac{1}{100e^{k}}$,
$V^{\pi^{*}}(s_{0})-V^{\tilde{\pi}_{k,m}}(s_{0})\leq\tilde{O}\left(\frac{1}{\sqrt{k}}\right),$
@@
G.1 Bounding the Bellman Error
This section is devoted to the proof of Lemma 4.3.
Theorem G.9 (Theorem 2 of Li et al. [2022]).
@@ G.1 Bounding the Bellman Error
$\|\hat{\mathcal{T}}_{k,m}\hat{Q}_{k,m}^{\mathrm{est}}-\hat{\mathcal{T}}_{k}\hat{Q}_{k}^{*}\|_{\infty}=\|\hat{Q}_{k,m}^{\mathrm{est}}-\hat{Q}_{k}^{*}\|_{\infty}\leq\epsilon_{k,m},$
@@
$m=\frac{c_{0}\cdot\tilde{r}\cdot t_{\text{cover}}}{(1-\gamma)^{5}\epsilon_{k,m}^{2}}\log^{2}\left(\frac{|\mathcal{S}||\mathcal{A}|}{\rho}\right)\log\left(\frac{1}{(1-\gamma)^{2}}\right),$ (30)
with probability at least $1-\rho$, for any $\rho\in(0,1)$. Here, $t_{\text{cover}}$ stands for the cover time, which is the time taken for the trajectory to visit all state-action pairs at least once. Formally,
@@
Lemma G.10.
If $T=\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}$, SUB-SAMPLE-MFQ: Learning runs in time $\tilde{O}(T|\mathcal{S}_{g}|^{2}|\mathcal{A}_{g}|^{2}|\mathcal{A}_{l}|^{2}|\mathcal{S}_{l}|^{2}k^{3.5+2|\mathcal{S}_{l}||\mathcal{A}_{l}|}\tilde{r})$, while accruing a Bellman noise $\epsilon_{k,m}\leq\tilde{O}(1/\sqrt{k})$ with probability at least $1-\frac{1}{100e^{k}}$, .
@@
We first prove that $\|\hat{Q}_{k}^{T}-\hat{Q}_{k}^{*}\|_{\infty}\leq\frac{1}{\sqrt{k}}$.
For this, it suffices to show $\gamma^{T}\frac{\tilde{r}}{1-\gamma}\leq\frac{1}{\sqrt{k}}\implies\gamma^{T}\leq\frac{1-\gamma}{\tilde{r}\sqrt{k}}$. Then, using $\gamma=1-(1-\gamma)\leq e^{-(1-\gamma)}$, it again suffices to show $e^{-(1-\gamma)T}\leq\frac{1-\gamma}{\tilde{r}\sqrt{k}}$. Taking logarithms, we have
$\exp(-T(1-\gamma))$ $\leq\frac{1-\gamma}{\tilde{r}\sqrt{k}}$
$-T(1-\gamma)$ $\leq\log\frac{1-\gamma}{\tilde{r}\sqrt{k}}$
$T$ $\geq\frac{1}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}$
Since $T=\frac{2}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}>\frac{1}{1-\gamma}\log\frac{\tilde{r}\sqrt{k}}{1-\gamma}$, the condition holds and $\|\hat{Q}_{k}^{T}-\hat{Q}_{k}^{*}\|_{\infty}\leq\frac{1}{\sqrt{k}}$. Then, rearranging Equation 30 and incorporating the convergence error of the $\hat{Q}_{k}$-function, one has that with probability at least $1-\rho$,
$\epsilon_{k,m}\leq\frac{1}{\sqrt{k}}+\frac{k\sqrt{2\tilde{r}t_{\text{cover}}}}{(1-\gamma)^{2.5}\sqrt{m}}\log\left(\frac{|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|}{\rho}\right)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$ (32)
@@
Substituting this in Equation 32 yields that with probability $1-\rho$,
$\epsilon_{k,m}\leq\frac{1}{\sqrt{k}}+\frac{k\sqrt{2\tilde{r}|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|}}}{(1-\gamma)^{2.5}\sqrt{m}}\log\left(\frac{|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|}{\rho}\right)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$ (33)
Therefore, letting $\rho=\frac{1}{100e^{k}}$, setting
$m=\frac{2\tilde{r}|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{S}_{l}||\mathcal{A}_{l}|k^{3.5+|\mathcal{S}_{l}||\mathcal{A}_{l}|}}{(1-\gamma)^{5}}\log\left({100|\mathcal{S}_{g}||\mathcal{A}_{g}||\mathcal{A}_{l}||\mathcal{S}_{l}|}{}\right)\log\left(\frac{1}{(1-\gamma)^{2}}\right)$ (34)
attains a Bellman error of $\epsilon_{k,m}\leq\tilde{O}(1/\sqrt{k})$ with probability at least $1-\frac{1}{100e^{k}}$. Finally, the runtime of our learning algorithm is
$O(mT|\mathcal{S}_{g}||\mathcal{S}_{l}||\mathcal{A}_{g}||\mathcal{A}_{l}|k^{|\mathcal{S}_{l}||\mathcal{A}_{l}|})=\tilde{O}(|\mathcal{S}_{g}|^{2}|\mathcal{A}_{g}|^{2}|\mathcal{A}_{l}|^{2}|\mathcal{S}_{l}|^{2}k^{3.5+2|\mathcal{S}_{l}||\mathcal{A}_{l}|}\tilde{r}),$
@@
Appendix H Generalization to Stochastic Rewards